OpenStudy (anonymous):

(sqrt)486 - (sqrt)24 + (sqrt) 6

OpenStudy (anonymous):

\[9\sqrt{6}-2\sqrt{6}+\sqrt{6}=8\sqrt{6}\]

OpenStudy (anonymous):

thank you